OpenStudy (dtan5457):

How many joules are needed to heat 48 grams of water from 10 degrees C to 96 degrees C? How many calories is this?

OpenStudy (dtan5457):

@ganeshie8

OpenStudy (dtan5457):

@wio

OpenStudy (dtan5457):

@Miracrown

OpenStudy (abb0t):

Use \(\sf \color{red}{q=mc \Delta T}\) q = heat m = mass c = heat capacity \(\Delta\)T = change in temperature (the difference)

OpenStudy (dtan5457):

which numbers represent what in this case?

OpenStudy (dtan5457):

@Kainui

OpenStudy (sweetburger):

what is the heat capacity of water?

OpenStudy (dtan5457):

i don't know. i wasn't here to learn this topic

OpenStudy (danjs):

@Abb0t is correct. Water has to absorb 4.184 Joules of heat for the temperature of one gram of water to increase 1 degree celsius (Specific Heat Capacity) \[c = \frac{ 4.184 J }{ g * C^{\circ }} \]

OpenStudy (abb0t):

As usual.

OpenStudy (danjs):

mass of 48 g, specific heat capacity above, change in temp of 86 degrees. To convert that to calories... the approximate amount of energy needed to raise the temperature of one gram of water by one degree Celsius at a pressure of one atmosphere... use 1 cal = about 4.2 Joules

OpenStudy (danjs):

ahah, mad chemist, of course as usual

OpenStudy (nechirwan):

specific heat capacity for water = 4.18 J/(g*degC) (how much energy is required per gram per change in degrees C) mass = 48g Change in temperature = 96-10 = 86 deg C energy required = mass * change in temperature * specific heat capacity = 48g * 86degC * 4.18 J/(g*degC) =17255.04 J 1 calories =4.184 joules or 1 joules =0.239005736 calories so 17255.04 *0.239005736=4124.05 calories